OpenStudy (anonymous):

Solve the equation. Show your work. 3x = 45

OpenStudy (anonymous):

divide 3 by both sides

OpenStudy (anonymous):

x=45/3

OpenStudy (epikrebel):

Basically \[\frac{3x}{3}=\frac{45}{3}\] x=15